Im trying to help my son, but I cant figure it out. Please Help.
Find two consecutive intergers so that the sum of the first and 3/4 of the second is 6.

Answer by checkley71(8403)

LET X & (X+1) BE THE 2 CONSECUTIVE NUMBERS.
X+.75(X+1)=6
X+.75X+.75=6
1.75X=6-.75
1.75X=5.25
X=5.25/1.75
X=3 THE FIRST INTEGER.
3+1=4 FOR THE OTHER INTEGER.
PROOF
3+.75*4=6
3+3=6
6=6